# Decouple and Track: Benchmarking and Improving Video Diffusion Transformers For Motion Transfer

### v1

@@ Abstract
The motion transfer task involves transferring motion from a source video to newly generated videos, requiring the model to decouple motion from appearance. Previous diffusion-based methods primarily rely on separate spatial and temporal attention mechanisms within 3D U-Net. In contrast, state-of-the-art video Diffusion Transformers (DiT) models use 3D full attention, which does not explicitly separate temporal and spatial information. Thus, the interaction between spatial and temporal dimensions makes decoupling motion and appearance more challenging for DiT models. In this paper, we propose DeT, a method that adapts DiT models to improve motion transfer ability. Our approach introduces a simple yet effective temporal kernel to smooth DiT features along the temporal dimension, facilitating the decoupling of foreground motion from background appearance. Meanwhile, the temporal kernel effectively captures temporal variations in DiT features, which are closely related to motion. Moreover, we introduce explicit supervision along dense trajectories in the latent feature space to further enhance motion consistency. Additionally, we present MTBench, a general and challenging benchmark for motion transfer. We also introduce a hybrid motion fidelity metric that considers both the global and local motion similarity. Therefore, our work provides a more comprehensive evaluation than previous works. Extensive experiments on MTBench demonstrate that DeT achieves the best trade-off between motion fidelity and edit fidelity.
@@ 1 Introduction
The advancement of foundation models for text-to-video generation has driven significant breakthroughs [17, 26, 50, 7, 18, 8]. Notably, Diffusion Transformers (DiT) video generation models have achieved remarkable results [37, 63, 59, 33]. However, generating complex motion remains a challenging, and relying solely on text prompts often fails to meet user expectations. To address this limitation, the motion transfer [60, 30, 23] seeks to generate videos that adhere to the motion in the source video under text control. This enables users to customize motion while retaining control over the foreground and background.
@@ 3.1 Motivations
To address this issue, we analyze the DiT feature $\mathcal{I}$ during the denoising process. As shown in Fig. 4, we observe that temporal inconsistencies in background features make distinguishing between foreground and background challenging. To separate foreground and background, an ideal approach is to smooth foreground and background features along the temporal dimension. Additionally, in order to accurately learn the motion pattern. We analyze the 3D attention map. The fig. 5 shows that strong diagonal attention scores appear only in adjacent frames. This suggests that modeling temporal changes in DiT features does not depend on the spatial receptive field and should instead focus on local temporal information. The reduction of spatial dimension also encourages the model to decouple the foreground appearance. Based on the above analysis, leveraging smoothness and temporal learning enables the decoupling of both foreground and background appearance, leading to successful motion transfer.
@@ 3.2 DeT
Table 1: Benchmark Comparison. We present a table summarizing the scale and difficulty of previous benchmarks.
Figure 7: Overview of MTBench. (a) illustrates the motion frequency distribution in evaluation prompts. (b) illustrates difficulty levels based on the number of trajectory clusters: 1-3 as easy, 4-6 as medium, and 7+ as hard.
@@
^$\hat{\mathcal{I}}_{xy,\,i}=\mathcal{I}^{\prime}_{xy,\,i}+\sum_{j=-\frac{k-1}{2}}^{\frac{k-1}{2}}\mathcal{I}^{\prime}_{xy,\,i+j}\times\mathcal{K}^{\text{t}}_{j}$ (1)
@@
~$\tilde{\mathcal{I}}=\mathcal{K}^{\text{t}}_{up}*\sigma\,(\mathcal{K}^{\text{t}}_{down}*\mathcal{I})+\text{Attention}(\mathcal{I},\mathcal{E}_{text})$ (2)
where $\mathcal{E}_{text}$ represents the embedding of the text prompt and * is the convolution operation. As a toy experiment on DiT, Fig. 4 (b) shows that the generated video follows the motion pattern of the source video while maintaining strong text controllability. Furthermore, Fig. 4 (c) demonstrates that our shared temporal kernel effectively enhances the model’s ability to distinguish between foreground and background.
@@
Specifically, we utilize CoTracker [25] to track the foreground in the source video, generating a set of trajectories $\mathcal{T}\in\mathbb{R}^{N\times T\times 2}$, where $N$ denotes the number of trajectories. Additionally, CoTracker provides the visibility of each point in the form of $\mathcal{V}\in\{0,1\}^{N\times T\times 1}$, where 1 indicates visibility, and 0 denotes invisibility. During the training process on source video $\mathcal{S}$, we calculate dense point tracking loss on the predicted latent feature ^$\hat{\mathcal{E}}(\mathcal{S})\in\mathbb{R}^{t\times h\times w\times c}$:x
^$\hat{\mathcal{E}}(\mathcal{S})$ $=\sqrt{\bar{\alpha}_{t}}\,\Bigl{(}\sqrt{\bar{\alpha}_{t}}\,\mathcal{E}(\mathcal{S})+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon\Bigr{)}-\sqrt{1-\bar{\alpha}_{t}}\,v$ (3)
@@
where =^$\mathcal{L}=\hat{\mathcal{E}}(\mathcal{S})\bigl{[}\mathcal{T}(t+1)\bigr{]}-\hat{\mathcal{E}}(\mathcal{S})\bigl{[}\mathcal{T}(t)\bigr{]}$. The final loss is a combination of denoising loss (DL) and dense point tracking loss (TL), weighted by $\lambda_{DL}$ and $\lambda_{TL}$:
@@ 3.3.1 Data Collection and Annotation
As shown in Tab. 1, previous benchmarks fall short in either scale (i.e., number of videos and prompts) or difficulty. To provide a more comprehensive evaluation of the motion transfer method, we construct a more general benchmark based on the DAVIS [40] and YouTube-VOS [57] datasets, which includes 100 high-quality videos with 500 evaluation prompts. We use the Qwen2.5-VL-7B [4] to annotate captions for the videos. Based on each caption, we generate five evaluation text prompts using the Qwen2.5-14B [58], swapping the foreground and background while keeping the verb unchanged. Additionally, we use the SAM [28] and CoTracker [25] to annotate masks and trajectories for the subjects in the videos. We then perform automatic K-means clustering of all trajectories based on the silhouette coefficient [12]. The motion in each video is categorized into easy, medium, and hard levels based on the number of clusters. We visualize the motion distribution in evaluation prompts and present examples of the difficulty division in Fig. 7 (a) and (b). We compared our benchmark with previous works in Tab. 1, showcasing the diverse range of motions it includes.
@@ 3.3.2 Metric
Hybrid motion fidelity (motion fidelity, for short). We introduce Fréchet distance $d_{F}$ [13] to compare the shape differences between trajectories $\mathcal{T}_{i},\,\mathcal{T}_{j}\in\mathbb{R}^{N\times T\times 2}$. It considers both the spatial arrangement and the order of points. This ensures that the transferred motion globally adheres to the source motion. We also use cosine similarity [60] to compare the local velocity direction $\Delta\mathcal{T}(t)=\mathcal{T}(t+1)-\mathcal{T}(t)$. Specifically, we compute motion fidelity score $\mathcal{M}$ through the formulation:
@@ 4 Experiment
Table 2: Quantitative results on MTBench. We denote EF: Edit Fidelity, TC: Temporal Consistency, MF: Motion Fidelity.
@@ 4.1 Experimental Setup
Baselines. We compare our method with state-of-the-art motion transfer methods, including MotionDirector [62], SMA [36], MOFT [56], and MotionClone [30]. Since previous motion transfer methods are based on 3D U-Net and cannot be directly applied to DiT, we implement DMT [60], MotionInversion [48], and DreamBooth [45] on DiT model to ensure a fair comparison.
@@ 4.2 Qualitative Results
As shown in Fig. 8, our method accurately transfers the motion from the source video without overfitting to the appearance. This enables flexible text control over both the foreground and background. Additionally, our method supports motion transfer across different categories, such as from a human to a panda or from a train to a boat. In Fig. 1 and Fig. 9, we compare our approach with previous motion transfer methods. U-Net-based methods [62, 36] struggle to maintain motion fidelity and exhibit weak text editing fidelity. We also adapt MotionInversion [48], DMT [60], and DreamBooth [45] to DiT [59]. However, we observe that Motion Inversion suffers from overfitting, failing to edit a rider into an astronaut or modify the background. DreamBooth and DMT fails to capture accurate motion patterns, causing originally curved motion trajectories to become linear. In contrast, our method accurately transfers motion patterns while allowing flexible text-based control over both the foreground and background appearance.
@@ 4.3 Quantitative Results
The quantitative results on MTBench are shown in Tab. 2. Our method with HunyuanVideo achieves the highest Motion Fidelity. While MOFT is slightly higher in edit fidelity, our method achieves the best balance between edit and motion fidelity. Additionally, we adapt MotionInversion, DreamBooth, and DMT to the DiT model, all of which underperform compared to our method across all metrics. Furthermore, in previous works, motion fidelity decreases as difficulty increases, while edit fidelity improves. This indicates that our benchmark difficulty division is reasonable. Notably, our method achieves the highest score on the medium difficulty level, which we attribute to its significant performance improvement in this level.
@@ 4.4 Ablation Study
Table 3: Ablation on different motion learning methods.
@@
Shared temporal kernel. We conduct the ablation study on the key design choices. Instead of using a shard temporal kernel, we consider alternative approaches through LoRA [21], Conv3D, and local attention [32]. As shown in Tab. 3, our method achieves the best performance in edit and motion fidelity metrics. Fig. 10 shows that shared temporal kernel achieve decoupling while learning the motion. We attribute our success to effective temporal smoothness and an attention map-based motion learning method. Additionally, we ablate the percentage of dropped layers, kernel size, and middle dimension in Tab. 4(a), Tab. 4(c), and Tab. 4(d), respectively. Quantitative results show that dropping 65% of layers, using a kernel size of 3, and setting the middle dimension to 128 achieve the best performance.
Dense point tracking loss. In Tab. 4(b), we ablate the weight of the dense point tracking loss and find that $\lambda_{TL}$ = 1e-1 achieves the best results. Figure 11 illustrates the cosine similarity between DiT features during training. The dense point tracking loss facilitates precise foreground tracking. Additionally, as shown in Fig. 12, we conduct a qualitative comparison for removing the dense point tracking loss. The green and red boxes highlight the effect of dense point tracking loss, which enhances the motion consistency.
@@ 9.2 DeT with HunyuanVideo
[…]

### v2

@@ Abstract
The motion transfer task aims to transfer motion from a source video to newly generated videos, requiring the model to decouple motion from appearance. Previous diffusion-based methods primarily rely on separate spatial and temporal attention mechanisms within the 3D U-Net. In contrast, state-of-the-art video Diffusion Transformers (DiT) models use 3D full attention, which does not explicitly separate temporal and spatial information. Thus, the interaction between spatial and temporal dimensions makes decoupling motion and appearance more challenging for DiT models. In this paper, we propose DeT, a method that adapts DiT models to improve motion transfer ability. Our approach introduces a simple yet effective temporal kernel to smooth DiT features along the temporal dimension, facilitating the decoupling of foreground motion from background appearance. Meanwhile, the temporal kernel effectively captures temporal variations in DiT features, which are closely related to motion. Moreover, we introduce explicit supervision along dense trajectories in the latent feature space to further enhance motion consistency. Additionally, we present MTBench, a general and challenging benchmark for motion transfer. We also introduce a hybrid motion fidelity metric that considers both the global and local motion similarity. Therefore, our work provides a more comprehensive evaluation than previous works. Extensive experiments on MTBench demonstrate that DeT achieves the best trade-off between motion fidelity and edit fidelity.
@@ 1 Introduction
The advancement of foundation models for text-to-video generation has driven significant breakthroughs [19, 29, 55, 7, 20, 8]. Notably, Diffusion Transformers (DiT) video generation models have achieved remarkable results [41, 68, 64, 36]. However, generating complex motion remains a challenging, and relying solely on text prompts often fails to meet user expectations. To address this limitation, the motion transfer [65, 33, 26, 44, 16, 17] seeks to generate videos that adhere to the motion in the source video under text control. This enables users to customize motion while retaining control over the foreground and background.
@@ 3 Method
Notation. We follow DDPM [23]: $t$ is the diffusion step, $\epsilon$ Gaussian noise, $\bar{\alpha}_{t}$ the cumulative noise schedule, and $\epsilon_{\theta}(\cdot)$ the noise predictor.
@@ 3.1 Motivations
To address this issue, we analyze the DiT feature $\mathcal{I}$ during the denoising process. As shown in Fig. 4, we observe that temporal inconsistencies in background features make distinguishing between foreground and background challenging. To separate foreground and background, an ideal approach is to smooth foreground and background features along the temporal dimension. Additionally, in order to accurately learn the motion pattern, we analyze the 3D attention map. The Fig. 5 shows that strong diagonal attention scores appear only in adjacent frames. This suggests that modeling temporal changes in DiT features does not depend on the spatial receptive field and should instead focus on local temporal information. The reduction of spatial dimension also encourages the model to decouple the foreground appearance. Based on the above analysis, leveraging smoothness and temporal learning enables the decoupling of both foreground and background appearance, leading to successful motion transfer.
@@ 3.2 DeT
Table 1: Benchmark Comparison. We present a table summarizing the scale and difficulty of previous benchmarks. The bottom row is the average number of clusters.
Figure 7: Overview of MTBench. (a) illustrates the motion frequency distribution in evaluation prompts. (b) illustrates difficulty levels based on the number of trajectory clusters: 1-3 as easy, 4-6 as medium, and 7+ as hard. The average number of clusters for the easy, medium, and hard levels are 2.7, 5.8, and 10.4, respectively.
@@
$\hat{\mathcal{I}}_{xy,\,i}=\mathcal{I}^{\prime}_{xy,\,i}+\sum_{j=-\frac{k-1}{2}}^{\frac{k-1}{2}}\mathcal{I}^{\prime}_{xy,\,i+j}\times\mathcal{K}^{\text{t}}_{j}$ (1)
@@
$\tilde{\mathcal{I}}=\mathcal{K}^{\text{t}}_{up}*\sigma\,(\mathcal{K}^{\text{t}}_{down}*\mathcal{I})+\text{Attention}(\mathcal{I},\mathcal{E}_{text})$ (2)
where $\mathcal{E}_{text}$ represents the embedding of the text prompt and * is the convolution operation. As a toy experiment on DiT, Fig. 3 (b) shows that the generated video follows the motion pattern of the source video while maintaining strong text controllability. Furthermore, Fig. 3 (c) demonstrates that our shared temporal kernel effectively enhances the model’s ability to distinguish between foreground and background DiT features.
@@
Specifically, we utilize CoTracker [28] to track the foreground in the source video, generating a set of trajectories $\mathcal{T}\in\mathbb{R}^{N\times T\times 2}$, where $N$ denotes the number of trajectories. Additionally, CoTracker provides the visibility of each point in the form of $\mathcal{V}\in\{0,1\}^{N\times T\times 1}$, where 1 indicates visibility, and 0 denotes invisibility. During the training process on source video $\mathcal{S}$, we calculate dense point tracking loss on the predicted latent feature $\hat{\mathcal{E}}(\mathcal{S})\in\mathbb{R}^{t\times h\times w\times c}$:
$\hat{\mathcal{E}}(\mathcal{S})$ $=\sqrt{\bar{\alpha}_{t}}\,\Bigl{(}\sqrt{\bar{\alpha}_{t}}\,\mathcal{E}(\mathcal{S})+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon\Bigr{)}-\sqrt{1-\bar{\alpha}_{t}}\,v$ (3)
@@
where $\mathcal{L}=\hat{\mathcal{E}}(\mathcal{S})\bigl{[}\mathcal{T}(t+1)\bigr{]}-\hat{\mathcal{E}}(\mathcal{S})\bigl{[}\mathcal{T}(t)\bigr{]}$. The final loss is a combination of denoising loss (DL) and dense point tracking loss (TL), weighted by $\lambda_{DL}$ and $\lambda_{TL}$:
@@ 3.3.1 Data Collection and Annotation
As shown in Tab. 1, existing benchmarks are limited in either scale (number of videos and prompts) or difficulty. To enable a more comprehensive evaluation of motion transfer, we construct a new benchmark based on DAVIS [45] and YouTube-VOS [62], consisting of 100 high-quality videos and 500 evaluation prompts. Captions are first generated using Qwen2.5-VL-7B [4], and five evaluation prompts per video are derived with Qwen2.5-14B [63] by swapping foreground and background while preserving the verb. Subject masks and trajectories are annotated with SAM [31] and CoTracker [28]. We then perform K-means clustering of trajectories based on the silhouette coefficient [12] to categorize motion into easy, medium, and hard levels by the number of clusters. Figure 7 (a) and (b) illustrate the motion distribution and difficulty division, while Tab. 1 highlights the broader motion diversity of our benchmark compared to prior works.
@@ 3.3.2 Metric
Hybrid motion fidelity (motion fidelity, for short). We introduce Fréchet distance $d_{F}$ [13, 37] to compare the shape differences between trajectories $\mathcal{T}_{i},\,\mathcal{T}_{j}\in\mathbb{R}^{N\times T\times 2}$. It considers both the spatial arrangement and the order of points. This ensures that the transferred motion globally adheres to the source motion. We also use cosine similarity [65] to compare the local velocity direction $\Delta\mathcal{T}(t)=\mathcal{T}(t+1)-\mathcal{T}(t)$. Specifically, we compute motion fidelity score $\mathcal{M}$ through the formulation:
@@ 4 Experiment
Table 2: Quantitative results on MTBench. EF: Edit Fidelity, TC: Temporal Consistency, MF: Motion Fidelity. For each column, the best score is shown in bold and the second‐best is underlined; the last two rows are not included in the ranking for a fair comparison.
@@ 4.1 Experimental Setup
Baselines. We compare our method against state-of-the-art motion transfer approaches, including MotionDirector [67], SMA [40], MOFT [61], and MotionClone [33]. Since prior methods rely on 3D U-Nets and cannot be directly applied to DiT, we additionally implement DMT [65], MotionInversion [53], and DreamBooth LoRA [50] on DiT for a fair comparison.
@@ 4.2 Qualitative Results
As illustrated in Fig. 8, our method faithfully transfers motion from the source video without overfitting to its appearance, enabling flexible text control over both foreground and background. It further supports cross‑category motion transfer, e.g., from a human to a panda or a train to a boat. Fig. 1 and 9 show that U‑Net‑based methods [67, 40] struggle with motion fidelity and exhibit weak text‑editing fidelity. When adapted to DiT, MotionInversion suffers from severe overfitting, failing to edit the rider into an astronaut or modify the background. DreamBooth and DMT fail to capture accurate motion, often straightening originally curved trajectories. In contrast, our method preserves fine‑grained motion patterns while supporting flexible, text‑driven edits.
@@ 4.3 Quantitative Results
Tab. 2 reports MTBench results. Our method, combined with HunyuanVideo, attains the highest motion fidelity and the most balanced trade‑off between edit and motion fidelity, with MOFT showing a marginal advantage in edit fidelity. Adaptations of MotionInversion, DreamBooth, and DMT to the DiT model consistently underperform across all metrics. As in prior studies, motion fidelity declines while edit fidelity improves with increasing difficulty, confirming the validity of our benchmark’s difficulty partitioning. Notably, our method achieves the top score at the medium level, driven by its substantial performance gains.
@@ 4.4 Ablation Study
Table 3: Ablation on different motion learning methods. The ablation experiments were run on a 30-video subset of MTBench, stratified to mirror the full benchmark’s distribution of difficulty levels and motion categories.
@@
Shared temporal kernel. We conduct the ablation study on the key design choices. Instead of using a shared temporal kernel, we consider alternative approaches through LoRA [24], Conv3D, and local attention [35]. As shown in Tab. 3, our method achieves the best performance in edit and motion fidelity metrics. Fig. 10 shows that shared temporal kernel achieve decoupling while learning the motion. We attribute our success to effective temporal smoothness and an attention map-based motion learning method. Additionally, we ablate the percentage of dropped layers, kernel size, and middle dimension in Tab. 4(a), Tab. 4(c), and Tab. 4(d), respectively. Quantitative results show that dropping 65% of layers, using a kernel size of 3, and setting the middle dimension to 128 achieve the best performance.
Dense point tracking loss. In Tab. 4(b), we ablate the weight of the dense point tracking loss and find that $\lambda_{TL}$ = 1e-1 achieves the best results. Fig. 11 illustrates the cosine similarity between DiT features during training. The dense point tracking loss facilitates precise foreground tracking. Additionally, as shown in Fig. 12, we conduct a qualitative comparison for removing the dense point tracking loss. The green and red boxes highlight the effect of dense point tracking loss, which enhances the motion consistency.
For additional ablation on the denoising loss weight $\lambda_{DL}$, please refer to the SM. Our qualitative and quantitative results show that reducing $\lambda_{DL}$ collapses background structures and increases edit fidelity, but simultaneously degrades motion fidelity.
@@ 7 More Qualitative Comparison Results
8 More quantitative Results
V2VBench. We further evaluate our approach on V2VBench. As shown in Tab. 5, DeT demonstrates a clear advantage in both motion and video-text alignment—key aspects of the motion transfer task.
User Study. We conduct a user preference evaluation. As shown in Tab. 10.4, DeT demonstrates a clear advantage across all assessed aspects.
@@ 9 Implementation Details
Table 5: V2VBench results (higher is better).
Method Motion Video Frames Object Semantic Video Frames Frames
Align $\uparrow$ Txt $\uparrow$ Txt $\uparrow$ Cons. $\uparrow$ Cons. $\uparrow$ Qual. $\uparrow$ Qual. $\uparrow$ Pick $\uparrow$
MotionDirector 3.09-3.09 $20.92$ 27.85 $0.94$ $0.95$ $0.62$ $4.98$ $0.26$
TokenFlow 1.57-1.57 $20.76$ $27.52$ $0.95$ $0.94$ 0.72 5.07 $0.25$
DeT (Ours) -0.83 31.50 $27.84$ 0.96 0.97 $0.63$ $4.99$ 0.29
@@ 10.1 Analysis for drop layers
Figure 21: Ablation of the denoising loss weight $\lambda_{DL}$.
10.2 Analysis for denoising loss weight
Our ablation on the denoising loss weight $\lambda_{DL}$ (Fig. 21, Tab. 10.4), shows that reducing $\lambda_{DL}$ collapses background structure and boosts edit fidelity, but simultaneously degrades motion fidelity.
@@ 10.3 DeT with HunyuanVideo
Dense point tracking loss. Fig. 6(b), we ablate the dense point tracking loss weight for DeT with HunyuanVideo. We find that $\lambda_{TL}=\text{1e-1}$ achieves the best results.
@@ 10.4 DeT with Step-Video-T2V
Table 8: Ablation on $\lambda_{DL}$.
$\lambda_{\text{DL}}$ EF TC MF
0.1 32.1 90.3 85.2
0.5 31.7 90.4 85.4
1.0 31.6 90.4 85.6
Table 9: User study results.
Method EF TC MF
MD 16.7% 20% 5%
MI 10% 23% 40%
DeT 73.3% 56% 55%